# Advancing Security with Digital Twins: A Comprehensive Survey

### v1

@@ III-B Domain-Based Application of Digital Twins
• Agriculture: Digital twin enables real-time monitoring of farms and greenhouses [32, 58, 59], allowing farmers to track key parameters like soil moisture, temperature, crop health, and weather patterns [60, 61]. Digital twin in agriculture supports resource optimization (e.g., water, energy) and early detection of pests or diseases. Manocha et al. proposed a smart irrigation framework using digital twins and IoT that processes soil, crop, and weather data to optimize water usage [62]. Digital twins have also been applied to greenhouse control using IoT and actuators [63, 64], and Moghadam et al. introduced a digital twin orchard for predicting crop stress and disease based on environmental simulations [65]. Chaux et al. developed a digital twin architecture to improve crop productivity in controlled environment agriculture by optimizing climate control and crop treatment strategies [66].
• Aviation: Digital twins are revolutionizing the aviation sector by enabling predictive maintenance, real-time structural health monitoring, and early anomaly detection, which are critical for enhancing safety, reliability, and improving operational efficiency. Given the high reliability demands of aerospace applications, Kapteyn et al. demonstrated the use of structural digital twins for unmanned aerial vehicle fleets to optimize operations through real-time sensor-driven simulations and mission planning [48], while Tuegel et al. introduced a high-fidelity aircraft digital twin updated with flight data to predict damage and lifespan using Bayesian learning [42]. Wang et al. combined fatigue mechanics with advanced filtering to forecast crack growth accurately [67], and Zakrajsek and Mall developed a physics-based digital twin for aircraft tires, using high-fidelity data and probabilistic models to predict wear and failure under varying touchdown conditions [68]. These studies underscore digital twins’ growing role in improving aerospace design, maintenance, and sustainability across the product lifecycle.
• Energy: Significant advancements have been made using digital twins to enhance the monitoring, optimization, and maintenance of power systems. Digital twins have been applied to renewable energy management, such as in wind and solar farms, to optimize energy output and reduce operational costs. GE [69], a pioneer in this field, showcased the use of digital twins in their digital wind farm to monitor turbine performance in real time and predict maintenance needs using sensor data. Li et al. proposed a data-driven digital twin approach for power electronic converters, using machine learning models (MLP, CNN, RNN) to monitor degradation in critical components like capacitors and MOSFETs, thereby improving reliability and efficiency through predictive maintenance [51]. Similarly, Wunderlich et al. developed real-time, embeddable digital twin models of power converters using NARX-ANNs, capable of performing fault detection, prognostics, health management, and risk assessment directly on embedded controllers [70]. In addition to performance optimization, digital twins have also shown potential for improving power system security, as demonstrated by Saad et al., who used IoT-based digital twins to detect and mitigate cyberattacks such as false data injection and denial of service, thereby enhancing the resilience of interconnected microgrids [71]. Further applications of digital twins in power systems can be found in [57, 72, 73, 74, 75].
• Healthcare: In healthcare, digital twins enable personalized and proactive treatment by digitally replicating patient physiology [76, 38]. Numerous smart healthcare applications have emerged, such as CloudDTH, which monitors, diagnoses, and predicts elderly health using wearable IoT devices [35], and a smart home healthcare that detects atrial fibrillation and falls using ECG and WiFi signals [77]. Thiong’o et al. proposed a DT to predict neurological complications in pediatric cancer patients by interfacing AI with patient-specific cancer profiles [78]. Another application is found in virtual medical training [79]. Ultimately, it enhances patient outcomes and overall quality of life.
@@
• Manufacturing: Digital twin is a key enabler of Industry 4.0, supporting real-time monitoring, prognostics, fault diagnosis, process optimization, and enhancing security. By leveraging digital twins, manufacturers gain agility, reduce waste, and enhance product quality. A prominent example is the Digital Twin Shopfloor (DTS) proposed by Tao et al., which integrates physical systems (PS), virtual systems (VS), service systems (SSS), and shop floor data (SDTD) to enable dynamic interaction for adaptive control and production planning. Real-time updates from the PS continuously refine the VS, enabling iterative optimization and correction [52, 28]. Similarly, Xu et al. introduced a Digital-Twin-Assisted Fault Diagnosis (DFDD) framework using deep transfer learning to overcome data scarcity in early production phases using synthetic data from virtual models to pretrain the system [36]. The pretrained model is then fine-tuned using real-world data to improve early fault detection, demonstrated in a case study on car body-side production. Additionally, Booyse et al. developed Deep Digital Twins (DDTs) trained solely on healthy operational data to detect asset degradation by identifying deviations and addressing the limitations of physics-based models in complex systems [37]. To secure Smart Factories, Salim et al. proposed a Blockchain-enabled Digital Twin framework that uses deep learning and smart contracts to detect botnets early by inspecting only packet headers without decrypting encrypted traffic, ensuring data integrity, privacy, and preventing bot spread [80]. Also, Becue et al. proposed a digital twin with cyber-range to assess attack impacts on avionics quality monitoring, enhance cobotic system resilience, and support iterative design of safety and security functions [81]. Collectively, these studies highlight digital twins’ capabilities in enabling accurate diagnostics, health monitoring, data privacy, and real-time decision-making.
• Smart Home: The integration of digital twin technology in smart homes represents a transformative advancement in home automation, integrating advanced technologies, such as IoT sensors, machine learning, actuators, and other interconnected devices, to create an intelligent ecosystem that enhances comfort, security, energy efficiency, and convenience for residents. For example, Liu et al. proposed a DT framework to improve the indoor safety of intelligent buildings. The system processes data collected by IoT sensors to automatically obtain the types and levels of danger in the building [82]. In [83], S.H. Khajavi et al. demonstrated the use of digital twins for safety and fire detection. Gopinath et al. introduced the re-design of smart homes using DT, integrating sensors with fans, front door locks, and lighting systems to enable remote control and real-time power usage monitoring. The integration of DT further allows engineers to test solutions virtually before physical maintenance, therefore minimizing unplanned downtime and improving system reliability [84]. Additionally, the implementation of DT in smart homes contributes to efficient usage of energy at home [85, 84].
@@
• Transportation: Kumar et al. introduced an AI-driven traffic congestion management system that integrates real-time data analytics, machine learning, blockchain, and digital twins. Central to this system is the Virtual Vehicle (VV) model, a digital replica of the driver and vehicle, which uses LSTM-based neural networks and edge/fog computing to predict driver behavior, analyze traffic patterns, and recommend optimized routes [86]. To ensure security, Arya et al. proposed a DT framework with machine learning to secure Vehicular ad-hoc networks by detecting malicious nodes and preventing DDoS attacks [87]. DTs are also advancing EV energy management and component health monitoring. Tesla, for instance, has adopted large-scale DTs by assigning each vehicle a unique virtual counterpart [88], enabling real-time sensor data processing for predictive maintenance, improved driving experiences, and over-the-air updates. Ye et al. proposed a DT-enhanced Q-learning energy management system that simulates driving forces and degradation dynamics to optimize energy use and extend battery life [89]. Venkatesan et al. proposed an intelligent digital twin (i-DT) framework for monitoring and predicting the health of electric motors in EVs using MATLAB/Simulink [90]. The system employs ANN and fuzzy logic to map operational data, such as travel distance, temperature, and magnetic flux, to critical outputs like lubricant schedules and remaining useful life. Their dual-layer approach combines in-vehicle monitoring with a cloud-based Remote Health Monitoring and Prognosis Center for real-time performance tracking, enhancing motor reliability and efficiency. In broader traffic applications, Kuši’c et al. developed the Geneva Motorway digital twin (DT-GM), which synchronizes real-time traffic data with SUMO-based microscopic simulations, updating every minute to dynamically model highway conditions [91].
@@ IV Digital Twin for Security
The DT framework presents itself as the ideal tool to enhance security in electronic devices due to its inherent abilities in detecting anomalies (e.g., vulnerabilities/flaws in design) from data, and the formulation of robust security strategies through playing what-if scenarios. The potential of DT in this domain has been widely acknowledged by security researchers, hence the need for this study [92, 93, 94, 18, 95, 96, 97]. In this section, we present recent studies on the exploration of DT for securing electronic systems. Figure 3 shows how DTs have been employed to address security concerns, including cyber-physical systems, counterfeit hardware, IoT devices, and information leakage, and is described in the following.
@@ IV-A2 Intrusion Detection
DTs are increasingly being used to enhance the security of CPS by tapping into their anomaly detection ability to detect intrusions in systems. Here, we review studies that utilize DTs to address intrusion detection and cyberattacks in CPS. Akbarian et al. proposed a digital twin-based intrusion detection system that leverages advanced techniques such as Kalman filter and support vector machine for attack detection and classification [101]. Balta et al. proposed a DT framework designed to detect cyberattacks in cyber-physical manufacturing systems by monitoring both controlled transient behaviors and anticipated process anomalies using run-time data [102]. El-Hajj used DT with an intrusion detection system to secure physical devices in CPS [103]. The framework effectively detected Hping3 attacks while also showing how different cyberattacks affect specific parts of a system. By understanding these effects, designers are able to create security defenses that are better suited to the system’s architecture and workload. Gehrmann and Gunnarsson introduced a DT-based security architecture to protect industrial automation and control systems (IACS) when opening up IACS low-level control functions and data exchange [104].
@@ IV-A3 Policy Generation
Traditionally, security policies are defined manually by domain experts to provide protection for an organization’s communication and computing infrastructures. As infrastructure update cycles accelerate and cyberattacks grow in complexity, manually defined policies struggle to keep pace with emerging threats. To address this challenge, Hammar and Stadler proposed digital twins as a solution for automating the development of effective security policies by simulating security scenarios in a controlled, virtual environment. Utilizing the data collected from simulations, reinforcement learning models are trained to automatically derive robust and adaptive security policies [97]. This approach not only enhances policy accuracy but also reduces human error and response time in rapidly evolving threat landscapes.
@@
IV-B Counterfeit Electronics
The increase in counterfeit electronic parts poses significant reliability and security threats to critical applications in defense, aircraft, vehicles, and medical devices due to their substandard quality, leading to major concerns for both government and industry [3, 105]. As most electronic devices are produced with limited oversight and pass through insecure supply chains, they are vulnerable to malicious compromise. For example, an adversary can tamper with the device and create a backdoor, which can be exploited for malicious purposes. When deployed in systems, the compromised devices can jeopardize the entire infrastructure. Note that counterfeit electronics are non-authentic electronics originating from the recycling of used components from discarded electronics, defective/out-of-spec parts, overproduced, cloned, forged documentation, tampered, and remarking of low-grade components to their higher grade counterpart (e.g., commercial grade components as military grade components).
@@ IV-B2 Supply Chain Management
With the rise of counterfeit electronics posing serious reliability and security risks to critical systems, DTs can offer a promising solution to enhance traceability and transparency across the supply chain. Leveraging data collected throughout the electronics lifecycle, DTs can provide insight into the root cause of component failures and support the detection of anomalies and unauthorized modifications. These capabilities can position DTs as a powerful tool for safeguarding electronic systems and ensuring the authenticity and integrity of devices from production to deployment.
@@ IV-C1 Data Privacy and Security
Low-powered devices often suffer from limited computational resources, memory, and power capacity, making them vulnerable to security breaches and privacy violations. To improve the privacy and security of the data being collected, Kumar et al. proposes a blockchain and deep learning-based framework to enhance security and data privacy in DT-empowered Industrial IoT networks by ensuring secure data transmission through smart contracts. The framework also detects attacks using a hybrid long short-term memory - sparse autoencoder and multi-head self-attention (MHSA)-based bidirectional gated recurrent unit intrusion detection model [106]. Similarly, in [107], the authors proposed the use of multiple digital twins for a single physical device to enhance data security in IoT devices. This use of multiple twins enables the separation of data among different virtual objects based on assigned tags and hence limits access to different data points to authorized users or applications, thereby preventing unauthorized data access.
@@ IV-D1 Side Channel Analysis
Side-channel attacks exploit the implementation flaws in cryptographic algorithms to expose secret keys by observing power consumption, electromagnetic emissions, and execution time during cryptographic operations. To address this issue, Yi et al. proposed a DT framework for side-channel resistant cryptographic devices by simulating the side channel information generated during the encryption process of cryptographic devices. By testing and analyzing the simulation model, the relevant side channel information and potential leakage points can be accurately captured and addressed. The authors tested different encrypted plaintexts, using a proposed measurement algorithm to accurately determine whether there is any side channel leakage [113].
@@ V Large Language Models in Digital Twin Frameworks
The preceding sections have explored how DTs revolutionize physical system modeling by creating dynamic virtual replicas that mirror real-world environments. Building upon this foundation, this section examines the integration of large language models (LLMs) with DT technology. As DTs evolve from static simulations to intelligent, predictive entities, LLMs emerge as powerful enablers that enhance their cognitive capabilities. This convergence creates systems that not only replicate physical counterparts but also understand, communicate, and reason about them in human-like ways. The following discussion explores the evolution of LLMs, their broad applications across domains, and their specific implementations within DT frameworks.

### v2

@@ III-B Domain-Based Application of Digital Twins
• Agriculture: In agriculture, digital twins enable real-time monitoring of farms and greenhouses by tracking soil moisture, temperature, crop health, and weather [33, 59, 60]. It allows resource optimization as demonstrated in Manocha et al. work, where DT was used to optimize water usage during irrigation, crop stress prediction [61], and climate control optimization for indoor agriculture [62].
@@
• Aviation: Digital twins advance aviation through predictive maintenance, health monitoring, and anomaly detection to improve safety and efficiency. In [49], Kapteyn et al. demonstrated the effectiveness of digital twins in mission planning and UAV fleet optimization. It also finds usefulness in aircraft damage prediction by leveraging flight data and Bayesian learning [43]. Similarly, Wang et al. proposed an approach for diagnostics and prognostics of damaged aircraft by combining fatigue mechanics and filtering theories [63]. Another application of digital is found in the work of Zakrajsek et al., which demonstrated the use of digital twins to predict wear in aircraft tires under different touchdown conditions.
• Energy: DTs find several applications in the energy domain, such as real-time monitoring, optimization, and predictive maintenance. A typical example in this domain is General Electric’s digital wind farm, which uses DTs for real-time turbine monitoring and predictive maintenance [64]. In [52, 65], the authors utilized DTs to detect degradation, fault, prognosis, and health management [65]. Saad et al. used IoT-based DTs to secure microgrids against cyberattacks, improving system resilience [66]. Broader DT applications in power systems are detailed in [58, 67, 68, 69, 70].
• Healthcare: In healthcare, digital twins support personalized treatment by replicating patient physiology [71]. Beyond personalized treatment, it also enables virtual medical training [72], health monitoring, and diagnosis, such as CloudDTH [36]. Another application of DT in healthcare is found in [73], detecting atrial fibrillation and falls using ECG and WiFi signal.
• Manufacturing: Digital twin is a key enabler of Industry 4.0, supporting real-time monitoring, fault diagnosis, optimization, and enhanced security. An example of this application is discussed in [53, 29], where the authors proposed a Digital Twin Shopfloor (DTS) which integrates physical, virtual, and service systems for adaptive control at the shop floor. Xu et al. introduced a DT-based fault diagnosis framework using synthetic and real data for early fault detection [37]. Booyse et al. developed Deep DTs to detect degradation from healthy data [38]. To enhance the security of systems, Salim et al. proposed a blockchain-enabled digital twin that uses smart contracts to detect botnets while preserving data privacy [74]. Becue et al. applied DTs to evaluate attack impacts and strengthen system resilience [75].
• Smart Home: DTs enhance smart homes by leveraging IoT sensors, ML, and actuators to improve comfort, security, and energy efficiency. Liu et al. proposed a DT framework for indoor safety using IoT sensor data [76], while Khajavi et al. applied DTs for safety and fire detection [77]. Gopinath et al. redesigned smart homes with DTs, enabling remote control of appliances and real-time energy monitoring [78]. DTs also support virtual testing and reduce downtime, boosting system reliability and energy efficiency [79, 78].
@@
• Transportation: Kumar et al. introduced an AI-driven traffic management system using digital twins, blockchain, LSTM models, and edge/fog computing to predict driver behavior and optimize routes [80]. In another work, Ye et al. used DT-enhanced Q-learning to simulate driving conditions and battery degradation, improving energy efficiency [81]. Venkatesan et al. proposed an intelligent DT for electric motor health monitoring using ANN and fuzzy logic, combining onboard diagnostics with cloud-based prognosis [82]. For traffic modeling, Kuši’c et al. developed the Geneva Motorway DT (DT-GM), which integrates real-time data with SUMO simulations and is updated every minute to reflect the dynamic highway conditions [83]. From a security perspective, Arya et al. developed a DT framework with machine learning to detect malicious nodes and mitigate DDoS attacks in vehicular networks [84]. Tesla also currently utilizes digital twins to enable predictive maintenance, over-the-air updates, and real-time analytics [85].
@@ IV Digital Twin for Security
The DT framework presents itself as the ideal tool to enhance security in electronic devices due to its inherent abilities in detecting anomalies (e.g., vulnerabilities/flaws in design) from data, and the formulation of robust security strategies through playing what-if scenarios. The potential of DT in this domain has been widely acknowledged by security researchers, hence the need for this study [86, 87, 88, 18, 89, 90, 91]. In this section, we present recent studies on the exploration of DT for securing electronic systems. Figure 3 shows how DTs have been employed to address security concerns, including cyber-physical systems, counterfeit electronics, IoT devices, and cryptographic systems, and is described in the following.
@@ IV-A2 Intrusion Detection
DTs are increasingly being used to enhance the security of CPS by tapping into their anomaly detection ability to detect intrusions in systems. Here, we review studies that utilize DTs to address intrusion detection and cyberattacks in CPS. Akbarian et al. proposed a digital twin-based intrusion detection system that leverages advanced techniques such as Kalman filter and support vector machine for attack detection and classification [95]. Balta et al. proposed a DT framework designed to detect cyberattacks in cyber-physical manufacturing systems by monitoring both controlled transient behaviors and anticipated process anomalies using run-time data [96]. El-Hajj used DT with an intrusion detection system to secure physical devices in CPS, effectively detecting Hping3 attacks while also showing how different cyberattacks affect specific parts of a system [97]. By understanding these effects, designers are able to create security defenses that are better suited to the system’s architecture and workload. Gehrmann and Gunnarsson introduced a DT-based security architecture to protect industrial automation and control systems (IACS) when opening up IACS low-level control functions and data exchange [98].
@@ IV-A3 Policy Generation
Traditionally, security policies are defined manually by domain experts to provide protection for an organization’s communication and computing infrastructures. As infrastructure update cycles accelerate and cyberattacks grow in complexity, manually defined policies struggle to keep pace with emerging threats. To address this challenge, Hammar and Stadler proposed DTs as a solution for automating the development of effective security policies by simulating security scenarios in a controlled, virtual environment. Utilizing the data collected from simulations, reinforcement learning models are trained to automatically derive robust and adaptive security policies [91]. This approach not only enhances policy accuracy but also reduces human error and response time in rapidly evolving threat landscapes.
@@
IV-B Supply Chain
The increase in counterfeit electronic parts poses significant reliability and security threats to the electronic supply chain supporting critical applications in defense, aircraft, vehicles, and medical devices due to their substandard quality, leading to major concerns for both government and industry [3, 99]. As most electronic devices are produced with limited oversight and pass through insecure supply chains, they are vulnerable to malicious compromise. For example, an adversary can tamper with the device and create a backdoor, which can be exploited for malicious purposes. When deployed in systems, the compromised devices can jeopardize the entire infrastructure. Note that counterfeit electronics are non-authentic electronics originating from the recycling of used components from discarded electronics, defective/out-of-spec parts, overproduced, cloned, forged documentation, tampered, and remarking of low-grade components to their higher-grade counterpart (e.g., commercial-grade components as military-grade components).
@@ IV-B2 Counterfeit Detection
With the rise of counterfeit electronics posing serious reliability and security risks to critical systems, DTs can offer a promising solution to enhance traceability and transparency across the supply chain. Using data collected throughout the electronics lifecycle, DTs can provide insight into the root cause of component failures and support the detection of anomalies and unauthorized modifications. These capabilities can position DTs as a powerful tool for safeguarding electronic systems and ensuring the authenticity and integrity of devices from production to deployment.
@@ IV-C1 Data Privacy and Security
Low-powered IoT devices are often constrained by limited computational resources, making them susceptible to security breaches and privacy violations. To address this, authors in [100] proposed a novel approach that employs multiple digital twins for a single physical device. By segmenting data across distinct virtual representations using assigned tags, this method enables fine-grained access control, ensuring that only authorized users or applications can access specific data points, thereby preventing unauthorized data access.
@@ IV-D1 Side Channel Analysis
Side-channel attacks exploit the implementation flaws in cryptographic algorithms to expose secret keys by observing power consumption, electromagnetic emissions, and execution time during cryptographic operations. To address this issue, Yi et al. proposed a DT framework for side-channel resistant cryptographic devices by simulating the side channel information generated during the encryption process of cryptographic devices. By testing and analyzing the simulation model, the relevant side channel information and potential leakage points are captured and addressed. The authors tested different encrypted plaintexts, using a proposed measurement algorithm to accurately determine whether there is any side channel leakage [107].
@@ IV-D2 Thermal Analysis
IV-E Comparative Assessment
Table II: A Summary Study of Digital Twins for Security.
Domains Scope Refs. Datasets [Availability] Accuracy Technologies TRL Limitations
CPS Security Analysis [93] Simulation data [N/P] N/A Cloud computing, IoT M Real-world validation
[94] Simulated network traffic between PLCs [P] N/A Sensors, Networking M Lacks performance evaluation, Cyber Threat integration
[86] TCP/IP network traffic [N/P] N/A Networking M Real-world validation, pipeline automation
Intrusion Detection [96] Run-time Data (temperature, motor counts) [P] N/A AI/ML, IoT M Detecting concurrent attacks and anomalies
[97] Simulation data (CPU, Memory, Network Usage)[A/R] 100% Data analytics M Real-world validation, Automated defense mechanism
[95] Simulation data [N/P] >83>83% Cloud computing M Real-world validation
Policy Gen. [91] Network traffic [N/P] N/A Networking, AI/ML M Generalizability to heterogeneous IT infrastructures
Supply Chain Counterfeit Detection [18] N/A N/A AI/ML L Data accessibility
[87] N/A N/A RFID, Blockchain, AI/ML L Empirical testing and evaluation
IoT Devices Data Privacy [100] Simulation data [N/P] 99% Networking, IoT M Interoperability across heterogeneous IoT platforms
Threat Detection [101] N/A N/A AI/ML, Blockchain, IoT L Lacks performance evaluation, hardware implementation
[102] N/A N/A AI/ML, Big data, IoT L Real-world validation
[103] Edge-IIoTset, ToN-IoT [P] >79>79% AI/ML, Cloud computing, IoT M High computational overhead
Security Manage- ment [104] Sensor data, Cyber threat intelligence data (NVD, CVE) [P] N/A Cloud computing, IoT M Limited scalability assessment
[105] Simulated network data [N/P] N/A IoT, Data analytics, Networking M Scalability, data availability
[106] Configuration and interaction data [N/P] N/A IoT, Blockchain M Real-world validation
Crypto-system SCA [107] Simulation data (power traces) [N/P], DPA contest v4 [P] N/A AI/ML M Generalization and scalability
Thermal Analy. [88] Real thermal data [N/P] N/A Semiconductor packaging M High computation
  • P: Published; A/R: Available on request; N/P: Not published; N/A: Not available; L: Low; M: Medium.
This section provides a comprehensive comparative assessment of existing DT-based security studies across various domains. Specifically, we analyze each study in terms of its performance metrics, enabling technologies, Technology Readiness Level (TRL), and dataset. Furthermore, we highlight limitations and open research challenges in this evolving field. The studies reviewed span several critical domains, each of which presents unique requirements and security concerns that influence the design and application of DT solutions.
Table II provides a detailed summary of recent research efforts that utilize digital twin technologies to address security challenges across a range of application domains. Column 1 specifies the domain in which each study is conducted (e.g., CPS, supply chain, IoT devices, and cryptosystem), while Column 2 outlines the specific scope or focus of the study within that domain (see Figure 3). Column 3 lists the corresponding references, pointing to the original research articles. Column 4 describes the types of datasets employed in these studies, which may include simulation-generated data, real-time operational data, or publicly available datasets such as those from the National Vulnerability Database (NVD) and Common Vulnerabilities and Exposures (CVEs). Column 5 presents the reported performance metrics, typically accuracy, where available (e.g., in [95], [97], etc.), and is marked as “N/A” when such data is not provided. Column 6 outlines the enabling technologies employed, including machine learning, cloud computing, IoT, and blockchain frameworks. Column 7 indicates the technology readiness level of each study, reflecting its degree of maturity: for simplicity, a low TRL corresponds to the conceptual research phase, medium indicates experimental validation, and high denotes deployment in real-world environments. Finally, Column 8 summarizes the major open challenges and limitations identified by each study, highlighting areas for future research and development in the field of DT-based security. While many of the reviewed studies demonstrate the promising potential of DT technology in enhancing security across domains such as IoT, CPS, and cryptographic systems, the majority remain at the experimental stage [93, 94, 86, 96, 97, 91, 95, 100, 103, 104, 105, 106, 107, 88]. A smaller subset of works, including [18, 101, 102, 87], remain in the conceptual phase with no reported real-world implementation or validation. Although applications in CPS and IoT have received comparatively more attention, including some hardware-based experimentation, these efforts are still constrained by challenges related to scalability, integration into existing infrastructures, and deployment in operational settings. Overcoming these limitations is essential for advancing DT-based security solutions from research prototypes to robust, industry-grade implementations.
@@ V Large Language Models in Digital Twin Frameworks
The preceding sections have explored how DTs revolutionize physical system modeling by creating dynamic virtual replicas that mirror real-world environments. Large language models (LLMs), with their natural language processing capabilities, have demonstrated their potential to enhance the cognitive abilities of DTs. This convergence creates systems that not only replicate physical counterparts but also understand, communicate, and reason about them in human-like ways. The following discussion explores the evolution of LLMs, their broad applications across domains, and their specific implementations within DT frameworks. The following discussion contains an overview of the integration of LLMs in DT frameworks and, especially, security. With an objective to analyze the existing studies and their limitations, we outline a thorough survey methodology and anchor the use cases of LLMs in distinct DT flows.